OpenStudy (anonymous):

Use the disk or the shell method to find the volume of the solid generated by revolving the region bounded by the graph of the equations about y = 6. y = 6/x^2 y = 0 x = 1 x = 3

OpenStudy (anonymous):

I used the washer method, and for some reason, I am continuing to do it wrong

OpenStudy (watchmath):

\(\int_1^3 \pi(6^2-(6-\frac{6}{x^2})^2)\,dx\)

OpenStudy (anonymous):

yes, that's what I started out with, using (top- bottom)

OpenStudy (watchmath):

ok, then how do you know that you continuing it wrong?

OpenStudy (anonymous):

\[\pi \int\limits_{1}^{3} 36 - ((36/x^4) - (72/x^2) +36) dx\]

OpenStudy (anonymous):

when I simplify, I am getting negative numbers, or no right answer at all

OpenStudy (anonymous):

\[\pi \int\limits_{1}^{3}((36/x4)−(72/x2))dx\]

OpenStudy (anonymous):

\[\pi \int\limits_{1}^{3}(-36/x^4) + (72/x^2) dx\]

OpenStudy (watchmath):

\(\pi\int_1^3 72x^{-2}-36x^{-4}\,dx\) \(\pi(-72x^{-1}+12x^{-3}\mid_1^3)\) Did you get that?

OpenStudy (anonymous):

no, I think maybe I did not integrate properly

OpenStudy (anonymous):

Everything you put up so far was right ilovemath

OpenStudy (anonymous):

I think the answer is 328/9 pi, so I will check it really quick

OpenStudy (anonymous):

Thanks for helping! The answer is correct!